# Automatic Outlier Rectification via Optimal Transport

### v1

@@ 5 Experimental Results
In this section, we demonstrate the effectiveness of the proposed statistically robust estimator through various simulated tasks: mean estimation and LAD regression.
5.1 Mean Estimation
Corruption Levels 20% 30% 40% 45% 49%
[…]
Ours, %Rectified 10.03% 10.12% 10.24% 10.35% 10.55%
Table 1: We compared our estimator with several standard mean estimation methods by evaluating the average loss on clean data points across various corruption levels. In our evaluation, we set the percent level of trimmed mean equal to the unknown ground truth corruption level. The hyperparameters for our estimator, namely $\delta=0.5$ and $r=0.5$, remained constant across all corrupted levels. The last row in the comparison table represents the percentage of all points rectified by our method.
We visualize the corrupted distribution and its rectified distribution when the corruption level is 45% in Figure 4. In this case, 55% of the data is drawn from the true distribution $\mathcal{N}(0,2)$ and 45% is drawn from $\mathcal{N}(25,2)$. Figure 4(a) displays the original sample of data from the contaminated distribution. The outlier points from the contaminating distribution are clearly visible in orange. Figure 4(b) shows the rectified distribution our estimator produces for $\delta=2.5$, in which the outlier points have been moved from their original values to their rectified values, which is much closer to the true mean of the clean distribution. Our estimator thus successfully identifies the majority of outliers and relocates them towards the center (our mean estimator), providing further support for our theoretical findings in the previous section. The sensitivity analysis of mean estimation with respect to $\delta$ is displayed in Figure 4(c). In this figure, the loss on the clean data is plotted for various values of $\delta$. We see that an approximate minimum occurs at $\delta=2$, with good performance within the range $\delta\in[0.5,2.5]$. This illustrates the relative insensitivity of our estimator to different values of $\delta$ for a wide and reasonable range. This range can be easily reached by hyperparameter tuning via cross-validation, as the function describing the performance of our estimator is approximately quasiconvex. Moreover, when $\delta=0$ or when $\delta$ is set so large that it can rectify all points, our estimator gracefully degrades to the loss of the median estimator, which is another favorable property. A further sensitivity analysis with respect to $r$ is given in Section C.1 in the Appendix. In this sensitivity analysis we find a similarly wide region of good performance for $r$ which improves significantly and almost uniformly on the more typical setting of $r=1$.
The evolution of the rectified distribution for mean estimation under concave and convex cost functions as $\delta$ increases is depicted in Appendices C.2 and C.3. These depictions show that our statistically robust estimator which applies the concave transport cost function acts in a stable and expected manner as $\delta$ increases. In particular, this depiction shows that our estimator moves points in the order of their distance from the estimate of the mean; that is, the budget is “used” on the worst outliers first. This results in an orderly procedure of rectification which is stable and predictable across nearby values of $\delta$, which is a favorable property if $\delta$ is being set sequentially or in cross-validation. In contrast, the estimator which applies the convex transport cost function moves all points of the clean and outlier distributions toward each other, which results in poor estimates of the true mean and an improper rectified distribution.
5.2 LAD Regression
As we discussed in the previous result, the theoretical analysis for the LAD estimator resembles that of the mean estimation task. Empirically, we also observe similar performance for the LAD estimator and show that our estimator correctly rectifies most of outliers to the fitted hyperplane under a range of choices for $\delta$. In order to further support the effectiveness of our estimator, we provide additional visualizations: Figure 5 contains a visualization of the lines of best fit produced by various estimations to the LAD regression problem, and displays the effect of different choices of $\delta$ on the line of best fit produced by our estimator. In Figure 5(a) our estimator with $\delta=1$ successfully produces a line of best fit which is closest to the uncontaminated distribution, while the other estimators (OLS, LAD, and Huber regressors) produce lines which are heavily affected by the contaminating distribution, showcasing the robustness of our estimator. Figure 5(b) and 5(c) show the rectified distribution produced by our estimator under $\delta=1$ (b) and $\delta=1.5$ (c). The suboptimal value of $\delta=1$ still rectifies many points from the contaminated distribution and produces a good line of best fit. Setting $\delta=1.5$ rectifies all of the points from the contaminated distribution and essentially recovers the true line of best fit. Importantly, this simulation shows that even improperly setting $\delta$ to the suboptimal value of $\delta=1$ produces a much better line than any of the other estimators in Figure 5(a). Appendix C.4 contains comparisons over different random trials and a sensitivity analysis with respect to $\delta$ and $r$, along with experimental results from our cross-validation approach.
6 Empirical Studies
In this section, we conduct empirical studies on real-world applications in fitting and predicting option implied volatility surfaces.
Options are financial instruments that grant the buyer the choice to buy or sell an asset at a predetermined price in the future. Those which grant buying privileges are referred to as call options, and those which grant sale privileges are referred to as put options. The predetermined price at which the underlying asset may be purchased or sold is known as the strike price. The option has an expiration date and time at which it expires. The remaining time until this point is referred to as the time to maturity. Options markets recognize various styles of exercising the right to buy or sell. In this empirical study, we only work with European-style options, in which exercise is only allowed at the expiration time. European options are broadly traded and analyzed in financial markets and are one of the most popular styles of options traded globally.
@@ 6 Empirical Studies
Option prices are generally impacted by the volatility of the underlying asset’s price process. The implied volatility (IV) is a measure derived from an option’s market price, reflecting the volatility expectation of the underlying asset’s price. The implied volatility surface (IVS) is the variation of implied volatility across options with different strike prices and times to maturity. The IVS is important in quantitative finance for risk assessment, hedging, pricing, and trading decisions (Cont & Da Fonseca, 2002; Daglish et al., 2007; Kamal & Gatheral, 2010; Gatheral, 2011). However, outliers in implied volatilities can distort the IVS and lead to inaccurate estimation. Consequently, there is a need for robust option implied volatility surface methods. Our approach handles this issue by estimating the surface in the presence of outliers. In this section, we perform two experiments which illustrate the efficacy of our statistically robust estimator for (a) estimating the IVS given options chains using a kernelized IVS estimator, and (b) estimating the IVS given options chains using a state-of-the-art deep learning IVS estimator.
6.1 Kernelized Volatility Surface Estimation
@@
Our data set consists of nearly 2,000 option chains containing end-of-day option prices from the US stock market, spanning the years 2019–2021. The options data comes from OptionMetrics via WRDS, a financial research data service available to major research universities. The chains chosen were identified as containing significant outliers by our industry partner, firm providing global financial data and analysis. We randomly divide each option chain into an 80% train set and a 20% test set. Using the train set, we estimate surfaces and evaluate their out-of-sample performance using mean absolute percentage error (MAPE) and the discretized surface gradient (defined as $\nabla\hat{S}$), a measure of surface smoothness. The MAPE evaluates the error of the options price surface when compared against the observed option implied volatilities. $\nabla\hat{S}$ evaluates the smoothness of the estimated surface. Further details of the data set can be found in Appendix C.8.
Losses.
These losses are chosen for their importance in finance and option trading and valuation. MAPE is a preferable error metric for volatility surfaces, as option chains contain options with implied volatilities which can differ in orders of magnitude, and MAPE weighs equally the contribution to error of options with such volatilities. The discretized surface gradient $\nabla\hat{S}$ measures surface smoothness of the estimated volatility surface by computing a discrete gradient across a grid of points in the (time to exporation, strike price) plane. Smooth surfaces are important for several reasons: (1) smoother surfaces allow more stable interpolation or extrapolation of implied volatility for options with market prices which are not directly observable; (2) smoother surfaces produce smaller adjustments for option hedging positions constructed from measures of the IVS, which ultimately lowers hedging transaction costs; (3) smoother surfaces are less likely to have internal arbitrages between options, which are implied by sharp discontinuities in the IVS, putting the surface more in line with established asset pricing theory; and (4) smoother surfaces produce more stable estimates of financial institutions’ derivative exposures, which are required to be consistent for financial regulatory and reporting requirements. These measures are standard in options IVS estimation, and further details on their definition can be found in Appendix C.8.
@@
We compare the kernelized statistically robust estimator developed in Theorem 2 to two benchmarks. The first is kernel smoothing (denoted “KS”), a well-established method for estimating the IVS (Aït-Sahalia & Lo, 1998). In kernel smoothing, each point on the IVS is constructed by a weighted local average of implied volatilities across nearby expiration dates, strike prices, and call/put type. The weights are given by a kernel matrix which depends on these three features of an option. The second benchmark is a two-step kernel smoothing method (denoted “2SKS”) which attempts to specially treat outlier points before estimating with the KS method. In this two-step method, options in an options chain are removed if their implied volatilities fall outside of the region $[q_{0.25}-1.5\cdot IQR,q_{0.75}+1.5\cdot IQR]$, where $q_{0.25}$ is the 25% quantile of the implied volatilities in the options chain, $q_{0.75}$ is defined similarly, and IQR is the interquartile range of the implied volatilities in the options chain. The surface is then estimated upon the remaining options via the KS method. This two-step “remove and fit” method is a well-known method for fitting models under outliers; the outlier identification procedure is commonly known as Tukey Fences (Hoaglin et al., 1986).
@@
Following the convention in (Aït-Sahalia & Lo, 1998), the $i$th option in a chain is featurized as the vector $\mathbb{R}^{3}\ni x_{i}^{\prime}=(\log\tau_{i},u(\Delta_{i}),\mathbb{I}_{\mathrm{call}(i)})$, where $\tau_{i}$ is the number of days to the option’s expiration date, $\Delta_{i}\in[-1,1]$ is a relative measure of price termed the Black-Scholes delta, $u(x)=x\mathbb{I}_{x\geq 0}+(1+x)\mathbb{I}_{x<0}$, and $\mathbb{I}_{\mathrm{call}(i)}$ is 1 if the option is a call option and 0 if the option is a put option. The goal is to fit the pair ($y_{i}^{\prime}$: the implied volatility of option $i$, $x_{i}^{\prime}$: the features) via a kernelized regression model. We choose a Gaussian-like kernel $K_{h}(x,x^{\prime})$ to measure the distance between options $x$ and $x^{\prime}$ as $K_{h}(x,x^{\prime})=\exp(-\|{(x-x^{\prime})}/{2h}\|^{2})$ where division of $x-x^{\prime}$ by $h$ is element-wise for a vector of bandwidths $h\in\mathbb{R}^{3}_{+}$. When the budget $\delta=0$, we want to solve $\min_{\theta\in\mathbb{R}^{n}}\sum_{i=1}^{n}v_{i}|y_{i}^{\prime}-\theta^{T}K_{h}^{i}|,$ where $v_{i}$ is the $i$-th entry of a vector of vegas for $\{x^{\prime}_{i}\}_{i=1}^{n}$ and $K_{h}^{i}$ is the $i$-th row of the kernel matrix. The implied volatilities are weighted by vega $v_{i}$ to improve surface stability as in (Mayhew, 1995). We conducted a comparison between our estimator and the benchmarks. We note that the KS method can be regarded as a standard kernelized least square approach (Hansen, 2022).
@@ Results.
Our approach improves upon the benchmark approach in both MAPE and $\nabla\hat{S}$ on the out-of-sample test sets, showcasing the importance of jointly estimating the rectified distribution and model parameters with our estimator. The results of our experiment are displayed in Table 2. We compare the KS and 2SKS benchmarks against our estimator with a fixed $\delta=0.01$ (chosen to correspond to a 1% change in volatility) and $\delta$ chosen by cross-validation, which is used to tune $\delta$ for each option chain. Details of the cross-validation procedure can be found in Appendix C.8.
Our estimator attains a mean MAPE of 0.225 across all chains, while the KS estimator and 2SKS estimator attain mean MAPEs of 0.294 and 0.236, respectively. The quantiles and mean of each estimator is reported in the upper section of Table 2 below. Our estimator thus improves upon the existing MAPE by 23% for the non-robust benchmark and 5% for the robust benchmark, illustrating the efficacy of our novel statistically robust estimator. The lower section of Table 2 reports the smoothness results of the estimated surfaces. Our estimator attains a mean $\nabla\hat{S}$ of 6.5 versus a mean of 20.2 for the KS estimator and 7.5 for the 2SKS estimator. This corresponds to improvements of 67% and 13%, respectively, demonstrating that outlier rectification is important for producing smooth surfaces which can be used in financial applications. [^1] It is notable to report that applying our estimator to an additional, completely out-of-sample test set collected after this paper was written allowed our industry partner to estimate and release to production 25% more surfaces than they had before when using their existing proprietary method.
@@
Table 2: This table reports the mean and quantiles of the distribution of MAPE and $\nabla\hat{S}$ across all option chains for the benchmark KS and 2SKS methods, and our estimator under $\delta=0.01$ and $\delta$ set via cross-validation. Our estimator achieves lower MAPE in many cases than the two benchmark approaches, and achieves much lower $\nabla\hat{S}$ than the benchmark estimators. Applying cross-validation generally improves smoothness, while maintaining similar MAPE. The improvement of our estimator over the benchmarks illustrates the importance of rectifying outliers for IVS estimation.
@@
We perform an additional experiment with the same dataset in Appendix C.7 which demonstrates the usefulness of our method for estimating an IVS for use on the trading day after the initial contaminated surface is observed. We find similar outperformance of our method versus the benchmark methods, with improvements in MAPE and $\nabla\hat{S}$. This experiment shows that our method is successful in interpolating as well as predicting the IVS in the presence of outliers.
6.2 Deep Learning Volatility Surface Estimation
In this section, we apply our statistically robust estimator to state-of-the-art deep learning prediction approaches for volatility modeling developed in Chataigner et al. (2020). We show that our estimator improves upon the performance of state of the art methods when faced with data containing outliers.
Background.
IVS estimation provides an approximate surface fit for the volatilities implied by option market prices. However, market participants often require surfaces which fit as closely as possible to the market-implied volatilities. [^2] For example, banks pricing some kinds of exotic options require such surfaces to avoid arbitrage opportunities in surfaces that their internal trading teams may exploit, as outlined in Hull & Basu (2022). Participants pricing options may moreover simply prefer to accept the assumption that the market prices are usually correct, and thus prefer closely fitting surfaces. One of the most popular methods used in mathematical finance to estimate surfaces obeying such a constraint under a reasonable stochastic process model for the evolution of asset prices is the local volatility or implied volatility function model introduced by Rubinstein (1994), Dupire et al. (1994), and Emanuel Derman & Zou (1996). The local volatility model assumptions imply that a function $\sigma(K,T)$ exists which gives an analytic expression for the volatility $\sigma$ at a given option strike price $K$ and option time to expiration $T$. This function is known as Dupire’s formula after its originator. Enforcing Dupire’s formula in the estimation of a surface results in a local volatility surface, which is of great use to financial market participants. However, estimating such a surface becomes significantly complicated if market data contains corruptions or unrealistic outliers which are unlikely to occur in the future. [^3] This may occur due to low liquidity, which allows market manipulation or adversarial trading to produce unrealistic prices. As human intervention for outlier rectification in a large number of IVS fitting tasks is unrealistic, an automatic method for rectifying outliers in models fitting local volatility surfaces is required.
Data.
We utilize the data set of Chataigner et al. (2020), which contains 17 (options chain, IVS) pairs taken from the German DAX index in August 2001. Options chains are captured from daily options market data, and IVSes are estimated for a grid of strikes and maturities via a trinomial tree calibrated by the method of Crépey (2002). In addition to Chataigner et al. (2020), this data set is also used in Crépey (2002) and Crépey (2004). In this data set’s usage in the literature, the authors train and test using the price surface, which is given by a nonlinear transformation of the IVS. Consequently, we also use the price surface for fitting; however, we report results after transforming back to the IVS. To corrupt the data set, a percentage of the options in each options chain are selected. We refer to this percentage, the corruption level, as $\epsilon$. The prices of these options $x$ are then corrupted by replacing them with the values $10x$.
Benchmark Method.
As a benchmark, we apply the state-of-the-art deep learning approach from Chataigner et al. (2020), the Dupire neural network. In this approach, a neural network estimates the price surface under the local volatility model assumptions encoded in Dupire’s formula, and additional no-arbitrage constraints which are useful for empirical applications. These conditions are enforced using different approaches: hard, split, and soft constraints. Hard constraints enforce the conditions by separating the network into subnetworks which have separate input layers for the variables involved in each of the conditions; conditions such as convexity and non-negativity are then enforced by projection steps during optimization and the proper choice of activation functions. Hard constraints thus learn a function which explicitly satisfies the conditions. Soft constraints enforce these conditions instead by penalizing violations of them at each of the training points observed from the market. Hybrid constraints split the network into subnetworks which have separate input layers for each of the constrained variables as in hard constraints, and utilize soft constraints for enforcing the Dupire and no-arbitrage conditions. The function learned only approximately satisfies the conditions in theory, but in practice the approximation is very good (see Chataigner et al. (2020) for details). Because these approaches estimate the price surface, we also estimate the price surface in this application; however, the local volatility surface is easily recovered from the price surface by a nonlinear transformation.
Robust Neural Network Estimation Problem.
To estimate price and volatility surfaces under the induced corruption, we estimated the benchmark approach using our statistically robust estimator. For each options chain, we tune $\delta$ using the cross-validation procedure outlined in the general procedure of Section 4.3.3 in the following way: for each day, we sample 80% of the training set without replacement as a cross-validation (CV) training set, and use the remaining 20% of the training set as a CV validation set. We then train our robust neural network estimator on the CV training set with $\delta\in\{10^{-3},10^{-2},...,10^{3}\}$. For each $\delta$, we then compute the MAPE on the CV validation set. We then select the $\delta$ with the lowest MAPE on CV validation set, and we use this $\delta$ to train our estimator on the entire training set. This produces our final estimated model.
Results.
@@
To maintain comparability with the benchmark, we test and report the same metrics reported by Chataigner et al. (2020), namely, the RMSE and MAPE for the surface estimation. The metrics for each day and model are repeated for three trials with different random weight initialization seeds. The results are displayed in Table 3 below. The first panel displays the RMSE and MAPE for each model using the different constraint techniques, averaged across all trials and corruption levels. The second panel displays the RMSE and MAPE for each corruption level averaged across all models. Our approach outperforms the baseline approach across all averages, and does so more clearly as the corruption level increases, despite the strong regularizing effect of the no-arbitrage constraints and enforcement of Dupire’s formula. Our improvement is present across all models, and is especially impactful for the most accurate model utilizing soft arbitrage constraints. For this model, the out-of-sample test error in terms of RMSE and MAPE are reduced by 33% and 34%, respectively. This experiment displays the efficacy of our estimator in a state-of-the-art deep learning model for option price surface modeling.
@@
Model
Hard Constraints
Hybrid Constraints
Soft Constraints
Dupire NN RMSE
0.125
0.140
0.044
Our RMSE
0.110
0.131
0.029
Dupire NN MAPE
0.343
0.546
0.111
Our MAPE
0.310
0.508
0.074
@@
Corruption Level
$\epsilon=20\%$
$\epsilon=30\%$
$\epsilon=40\%$
Dupire NN RMSE
0.077
0.091
0.168
Our RMSE
0.075
0.084
0.141
Dupire NN MAPE
0.061
0.070
0.115
Our MAPE
0.060
0.066
0.097
Table 3: Mean RMSE and MAPE for implied volatility estimation across all samples for the benchmark deep learning approach and the deep learning approach estimated using our statistically robust estimator, split by constraint method (top Panel A) and corruption level (bottom Panel B). Our estimator achieves significantly lower RMSE and MAPE in estimating the implied volatility surface. Results for models are averaged across $\epsilon\in\{0.2,0.3,0.4\}$, and results for corruption levels are averaged across models using the hard, hybrid, and soft constraint methods.

### v2

@@ 5 Experimental Results
In this section, we demonstrate the effectiveness of the proposed statistically robust estimator through various tasks: mean estimation, LAD regression, and two applications to volatility surface modeling.
5.1 Mean Estimation and LAD Regression Experiments
We perform simulation experiments with mean estimation and least absolute deviation linear regression for our estimator to illustrate its efficacy. For mean estimation, we compare our estimator with the mean, median, and trimmed mean estimators. We observe that our estimator outperforms all other methods, despite providing the oracle corruption level $\varepsilon$ to the trimmed mean estimator. For LAD regression, we compare our estimator with the OLS, LAD, and Huber regression estimators. We find that, again, our estimator outperforms all other methods. A complete overview of these two experiments can be found in Appendix E.
5.2 Options Volatility Experiments
@@
In this section, we conduct empirical studies on real-world applications in fitting and predicting option implied volatility surfaces. Options are financial instruments allowing buyers the right to buy (call options) or sell (put options) an asset at a predetermined price (strike price) by a specified expiration date. In this study, we focus on European-style options, which can only be exercised at expiration. Option prices are influenced by the volatility of the underlying asset’s price. Implied volatility (IV), derived from an option’s market price, indicates the market’s volatility expectations. The implied volatility surface (IVS) represents the variation of IV across different strike prices and times to maturity. Accurate IVS modeling is crucial for risk assessment, hedging, pricing, and trading decisions. However, outliers in IV can distort the IVS, necessitating robust estimation methods. Our approach addresses this by estimating the IVS in the presence of outliers. We conduct two experiments to demonstrate the effectiveness of our statistically robust estimator: (a) using a kernelized IVS estimator and (b) using a state-of-the-art deep learning IVS estimator.
5.2.1 Kernelized Volatility Surface Estimation
@@
Our data set comprises nearly 2,000 option chains containing daily US stock option prices from 2019–2021. The options data is sourced from WRDS, a financial research service available to universities. The option chains were identified as containing significant outliers by our industry partner, a firm providing global financial data and analysis. We were blinded to this choice. We randomly draw a training set and test set from each chain and estimate surfaces. We assess the surfaces’ out-of-sample performance using mean absolute percentage error (MAPE) and the discretized surface gradient (defined as $\nabla\hat{S}$). MAPE evaluates the error of the surface versus observed implied volatilities. $\nabla\hat{S}$ evaluates the smoothness of the surface. Further details can be found in Appendix G.2.
@@
We compare our estimator developed in Theorem 4 to two benchmarks. The first is kernel smoothing (denoted “KS”), a well-established method for estimating the IVS (Aït-Sahalia and Lo, 1998). In KS, each point on the IVS is constructed by a weighted local average of implied volatilities across nearby expiration dates, strike prices, and call/put type. The weights are given by a kernel matrix which depends on these three features of an option (see next paragraph). The second benchmark is a two-step “remove and fit” kernel smoothing method (denoted “2SKS”) which first attempts to remove outliers via Tukey fences (Hoaglin et al., 1986) before applying the KS method.
@@
We now describe the kernel and our estimator. Following the convention in (Aït-Sahalia and Lo, 1998), the $i$th option in a chain is featurized as the vector $\mathbb{R}^{3}\ni x_{i}^{\prime}=(\log\tau_{i},u(\Delta_{i}),\mathbb{I}_{\mathrm{call}(i)})$, where $\tau_{i}$ is the number of days to the option’s expiration date, $\Delta_{i}\in[-1,1]$ is a relative measure of price termed the Black-Scholes delta, $u(x)=x\mathbb{I}_{x\geqslant 0}+(1+x)\mathbb{I}_{x<0}$, and $\mathbb{I}_{\mathrm{call}(i)}$ is 1 if the option is a call option and 0 if the option is a put option. The goal is to fit the pair ($y_{i}^{\prime}$: the implied volatility of option $i$, $x_{i}^{\prime}$: the features) via a kernelized regression model. We choose a Gaussian-like kernel $K_{h}(x,x^{\prime})$ to measure the distance between options $x$ and $x^{\prime}$ as $K_{h}(x,x^{\prime})=\exp(-\|{(x-x^{\prime})}/{2h}\|^{2})$ where division of $x-x^{\prime}$ by $h$ is element-wise for a vector of bandwidths $h\in\mathbb{R}^{3}_{+}$. When the budget $\delta=0$, we want to solve $\min_{\theta\in\mathbb{R}^{n}}\sum_{i=1}^{n}v_{i}|y_{i}^{\prime}-\theta^{T}K_{h}^{i}|,$ where $v_{i}$ is the $i$-th entry of a vector of vegas for $\{x^{\prime}_{i}\}_{i=1}^{n}$ and $K_{h}^{i}$ is the $i$-th row of the kernel matrix. The implied volatilities are weighted by vega $v_{i}$ to improve surface stability as in (Mayhew, 1995). We conducted a comparison between our estimator and the benchmarks. We note that the KS method can be regarded as a standard kernelized least square approach (Hansen, 2022).
@@ Results.
Our approach improves upon the benchmark approach in both MAPE and $\nabla\hat{S}$ on the out-of-sample test sets, showcasing the importance of jointly estimating the rectified distribution and model parameters with our estimator. The results of our experiment are displayed in Table 1. We compare the KS and 2SKS benchmarks against our estimator with both fixed $\delta=0.01$ (chosen to correspond to a 1% change in volatility) and $\delta$ chosen by cross-validation. More details can be found in Appendix G.2. Our estimator outperforms others with a mean MAPE of 0.225, compared to 0.294 for KS and 0.236 for 2SKS, showing a 23% and 5% improvement respectively. For surface smoothness, our estimator achieves a mean DSG of 6.5, significantly better than 20.2 for KS and 7.5 for 2SKS, indicating 67% and 13% improvements. Notably, our industry partner applied our estimator to an unseen test set collected after this paper was written and was able to release to production 25% more surfaces than they had before when using their existing proprietary method.
Table 1: Results of our experiment with kernelized IVS estimation.
@@
Remark 5.1.
@@
We perform an additional experiment with the same dataset in Appendix G.1 which demonstrates the usefulness of our method for estimating an IVS for use on the trading day after the initial contaminated surface is observed. We find similar outperformance of our method versus the benchmark methods.
5.2.2 Deep Learning Volatility Surface Estimation
In this section, we apply our statistically robust estimator to state-of-the-art deep learning prediction approaches for modeling volatility developed in Chataigner et al. (2020). In this work, deep networks are used to estimate local volatility surfaces. The local volatility or implied volatility function model introduced by Rubinstein (1994), Dupire et al. (1994), and Emanuel Derman and Zou (1996) is a surface which allows for as close of a fit as possible to market-observed implied volatilies without arbitrage, which is of interest to many market participants (smoothing methods, in contrast, do not make this guarantee). Further background and details for this section are available in Appendix G.3.
We select the data set from Chataigner et al. (2020) consisting of (options chain, surface) pairs from the German DAX index. We first contaminate the chains by replacing an $\varepsilon$ fraction of each price $p$ with $10p$. We then test our estimated surface against the true surface. To estimate price and volatility surfaces under data corruption, we applied our statistically robust estimator to the benchmark approach. We use the Dupire neural network of Chataigner et al. (2020) as a benchmark, which estimates the surface under local volatility model assumptions and no-arbitrage constraints. This method enforces these conditions on the surface using hard, soft, and hybrid hard/soft constraints.
@@ 5.2.2 Deep Learning Volatility Surface Estimation
We evaluate our robust estimator using the same metrics as Chataigner et al. (2020), RMSE and MAPE, repeated over three trials with different random seeds. Our approach outperforms the baseline approach across all averages, and does so more clearly as the corruption level increases, despite the strong regularizing effect of the no-arbitrage constraints and enforcement of Dupire’s formula. Our improvement is most impactful for the most accurate model utilizing soft arbitrage constraints. For this model, the test set RMSE and MAPE are reduced by 33% and 34%. This experiment displays the efficacy of our estimator in a state-of-the-art deep learning approach to volatility modeling.
Table 2: Results of our experiment with deep learning surface estimation.
@@
Model Hard Constraints Hybrid Constraints Soft Constraints
Dupire NN RMSE 0.125 0.140 0.044
Our RMSE 0.110 0.131 0.029
Dupire NN MAPE 0.343 0.546 0.111
Our MAPE 0.310 0.508 0.074
@@
Corruption Level $\varepsilon=20\%$ $\varepsilon=30\%$ $\varepsilon=40\%$
Dupire NN RMSE 0.077 0.091 0.168
Our RMSE 0.075 0.084 0.141
Dupire NN MAPE 0.061 0.070 0.115
Our MAPE 0.060 0.066 0.097